# Sampling Theory for Function Approximation with Numerical Redundancy

### v1

@@ 5.2.2 Uniformly random sampling
6 Implicit regularization in numerical orthogonalization
It might be tempting to avoid working with numerically redundant spanning sets by simply orthogonalizing them. Indeed, several computational methods work with a numerically orthogonalized basis, for example [40, 23]. How do these methods compare to the regularized least squares method analysed in this paper? Are they substantially different? In this section, we show that regularization is implicitly present in straightforward numerical orthogonalization techniques. We aim to convey the main ideas, though each pointer provided in this section deserves further exploration. Note that Theorem 3 also suggests that it is impossible to orthogonalize a set numerically without suffering from the effects associated with finite precision.
6.1 Three orthogonalization strategies
For simplicity, we assume that a set of linearly independent functions $\{\eta_{1},\dots,\eta_{n}\}$ is given, which spans the approximation space $V_{n}$ with $\dim(V_{n})=n$. One straightforward way of orthogonalizing consists of discretizing these functions on a fine grid $\{\mathbf{t}_{j}\}_{j=1}^{l}$ and computing a factorization
This basis is orthogonal with respect to the discrete inner product defined by $\langle f,g\rangle_{l}=(1/l)\sum_{j=1}^{l}f(\mathbf{t}_{j})\overline{g(\mathbf{t}_{j})}$. If the grid $\{\mathbf{t}_{j}\}_{j=1}^{l}$ is sufficiently dense, the functions are also (approximately) orthonormal with respect to the inner product associated with the Hilbert space $H$. The basis $\{\widetilde{\eta}_{j}\}_{j=1}^{n}$ can then be used as a spanning set $\Phi_{n}=\{\phi_{i,n}\}_{i=1}^{n}$ (1) for least squares approximation.
Importantly, in the case of a numerically redundant set of functions $\{\eta_{1},\dots,\eta_{n}\}$, the matrices $A$ and $R$ are heavily ill-conditioned. More specifically, their singular value profiles mimick that of the synthesis operator associated with $\{\eta_{1},\dots,\eta_{n}\}$, having very small yet nonzero singular values. Can (67) even be computed with high accuracy in such cases? As an example, we consider the QR decomposition as a factorization method for (66), such that $R$ is upper-triangular. The accuracy of solving triangular systems of equations is analysed extensively in [33], which shows that the solution can be surprisingly accurate despite ill-conditioning. The key insight is that the forward error depends on the Skeel condition number [45]:
[…]
In this case, we do not have to solve an ill-conditioned linear system; however, we can only access the basis $\{\widetilde{\eta}_{1},\dots,\widetilde{\eta}_{n}\}$ on the grid $\{\mathbf{t}_{j}\}_{j=1}^{l}$. For an algorithm implemented in finite precision, it can only be guaranteed that the computed matrices $\hat{Q}$ and $\hat{R}$ factor the matrix $A$ with a small backward error. As an example, we consider the Householder QR algorithm, which is analysed in [34, section 19.3]. Following [34, p.361], one has
where $C$ is a constant that depends on the dimensions of the problem and the working precision $\operatorname{\epsilon_{\text{mach}}}$. Because of the small singular values of $A$, the range of $A+\Delta A$, and consequently the range of $\hat{Q}$, can differ significantly from that of $A$. A toy example confirming this statement is given in [47, p. 263], and a study of the angle between the subspaces $\operatorname{R}(A)$ and $\operatorname{R}(A+\Delta A)$ is given in [25]. In conclusion, there are no guarantees that the computed basis functions span the complete approximation space $V_{n}$. Observe the similarity between (69) and, for example, TSVD regularization: the range of $\hat{Q}$ contains the singular vectors of $A$ associated with large singular values $\sigma_{i}\gg\operatorname{\epsilon_{\text{mach}}}$, yet the directions associated to small singular values can be perturbed or even absent. This implies that the best approximation still cannot be computed.
Finally, there exist algorithms that do succeed in computing an orthonormal basis for the whole approximation space $V_{n}$, seemingly starting from a numerically redundant set. These methods typically use some additional analytic knowledge. One example is the Stieltjes procedure for generating orthogonal polynomials [27], which recently gained popularity for numerical polynomial approximations in the form of the Vandermonde with Arnoldi (VwA) algorithm [16]. The algorithm is built on the simple principle that multiplying a polynomial by $x$ increases its degree. The orthogonalization procedure can be summarized by the following Matlab code, using Chebfun [24]:
T_v = chebfun(’1’); x = chebfun(’x’); T_u = T_v/norm(T_v);
for i = 2:n
% new Arnoldi iteration = multiplication by x
v = x*T_u(:,end);
T_v = [T_v v];
% orthogonalize against previous functions
v = v - T_u*pinv(T_u)*v;
T_u = [T_u v/norm(v)];
end
Most importantly, the algorithm does not actually orthogonalize monomials. Instead, new “Arnoldi” basis functions, stored in T_v, are iteratively created to be better suited for orthogonalization. Indeed, the algorithm does not orthogonalize $x^{i}$, but $x$ times the previous orthonormal basis function. On Figure 8, the first eight of these Arnoldi basis functions are plotted, along with the singular value profile of the synthesis operator associated with this basis for $n=100$. Compare this to Figure 3(a) and 3(b). The synthesis operator associated with the Arnoldi basis is much better conditioned than the synthesis operator of the monomials. This phenomenon is analysed for a variety of domains in [48]. We remark that the VwA algorithm performs discrete orthogonalization. Furthermore, the algorithm can be extended to other sets of functions provided they have similar Vandermonde-like system matrices $A$ (66), as illustrated in the example below.
(a) first eight basis functions
(b) singular value profile of $\mathcal{T}_{n}$ for $n=100$
Figure 8: The Arnoldi basis associated with Stieltjes orthogonalization / the Vandermonde with Arnoldi algorithm along with the singular value profile of the associated synthesis operator. This set of functions has well behaved frame bounds, i.e., $B_{n}/A_{n}$ is small, in contrast to the monomials displayed in Figure 3(a) and 3(b). Hence, the set is not numerically redundant and can be numerically orthogonalized without suffering from the effects of finite precision.
6.2 A numerical example
As an illustration, the different methods mentioned above are compared numerically. To this end, we reproduce [16, Example 3]. More specifically, we approximate the function $f(x)=1/(10-9x)$ on $[-1,1]$ using a Fourier extension on $[-2,2]$, i.e.,
$f(x)\approx\sum_{k=-n}^{n}c_{k}\exp(ik\pi x/2).$
Since $f$ is real, this can be simplified to
On [16, Fig. 5.1], a comparison is shown between the direct computation of a least squares fit using Matlab’s backslash and the Vandermonde with Arnoldi (VwA) algorithm. The computations are based on function evaluations in 1000 Chebyshev points. Note that VwA is applicable in this setting, since the discrete least squares matrix has Vandermonde structure as described in [16, Example 3]. It is also important to remark that Matlab’s backslash implicitly computes a regularized least squares approximation due to the numerical rank-deficiency of the discrete least squares matrix.
We add some extra methods to the comparison: approximation using a numerically orthogonalized basis obtained via a QR decomposition as described above, and a truncated singular value (TSVD) approximation. For implementational details, we refer to [16] and [31]. The results are presented in Figure 9. The VwA algorithm achieves exponential convergence, which is to be expected in $\operatorname{span}(\Phi_{n})$ [35]. On the other hand, the accuracy of both the regularized and the numerically orthogonalized approximations decreases compared to the theoretical convergence behaviour around $n=20$. This example confirms that straightforward numerical orthogonalization techniques implicitly regularize the approximation problem.
Figure 10 shows the singular value profile of the synthesis operator associated with $\Phi_{n}$ for $n=10$ (left), $n=20$ (middle), and $n=30$ (right), computed in double precision using Chebfun [24]. Without rounding errors, the singular values would proceed to decrease exponentially toward zero as $n$ goes to infinity. From the frame bounds $A_{n}$ and $B_{n}$ of $\Phi_{n}$, one can infer that $\Phi_{n}$ becomes numerically redundant around $n=20$. As shown in Figure 9, this is the exact point where the influence of regularization on the convergence behaviour becomes significant.
Perpendicular to the question of how to obtain an orthonormal basis, one might wonder what the advantages and disadvantages are of approximation with an orthonormalized spanning set (for example, using the VwA algorithm) compared to regularized approximation with a redundant spanning set. Most importantly, as illustrated in Figure 9, regularization typically slows down convergence. On the other hand, orthogonalizing often throws away exploitable structure. More specifically, the discrete least squares matrix associated with $\Phi_{n}$ in the example above is a submatrix of a Fourier matrix. This structure can be exploited to design efficient least squares methods, see for example the AZ algorithm for frames approximations [20, 30], or to allow for cheaper evaluation of the approximation. Furthermore, regularization also lowers the number of measurements that are needed for accurate approximation, which is exactly the focus of this paper.
@@ Appendix B Proof of Theorem 10
$k_{n}^{\epsilon}(\mathbf{x})=z(\mathbf{x})(G_{n}+\epsilon^{2}I)^{-1}z(\mathbf{x})^{*},$
where $z(\mathbf{x})=\begin{bmatrix}\phi_{1}(\mathbf{x})&\dots&\phi_{n^{\prime}}(\mathbf{x})\end{bmatrix}\in\mathbb{C}^{1\times n^{\prime}}$ and $G_{n}$ represents $\mathcal{T}_{n}^{*}\mathcal{T}_{n}$.
@@
$G_{n}$ is positive semidefinite such that there exists an eigenvalue decomposition of the form $V\Sigma^{2}V^{*}$, where $V\in\mathbb{C}^{n^{\prime}\times n^{\prime}}$ is unitary and $\Sigma\in\mathbb{C}^{n^{\prime}\times n^{\prime}}$ is diagonal. Moreover, this implies that the synthesis operator $\mathcal{T}_{n}$ associated with $\Phi_{n}$ has a singular value decomposition of the form $\sum_{i=1}^{n}\sigma_{i}u_{i}\mathbf{v}_{i}^{*}$ where $\{u_{i}\}_{i=1}^{n}$ is an orthonormal basis for $V_{n}$, $\mathbf{v}_{i}$ are the first $n$ columns of $V$ and $\sigma_{i}=(\Sigma)_{i,i}$ for $i=1\dots n$. Therefore, using $z(\mathbf{x})=\sum_{i=1}^{n}\sigma_{i}u_{i}(\mathbf{x})\mathbf{v}_{i}^{*}$, we obtain
$z(\mathbf{x})(G_{n}+\epsilon^{2}I)^{-1}z(\mathbf{x})^{*}$ $=\left(\sum_{i=1}^{n}\sigma_{i}u_{i}(\mathbf{x})\mathbf{v}_{i}^{*}\right)V(\Sigma^{2}+\epsilon^{2}I)^{-1}V^{*}\left(\sum_{i=1}^{n}\sigma_{i}\overline{u_{i}(\mathbf{x})}\mathbf{v}_{i}\right)$
@@
The inequality on the left is exactly the $\Delta$-spectral approximation introduced in [10, eq. (2)] with $\Delta=1/2$. This allows us to reuse the method of proof of [10, Lemma 6]. Analogously, using the eigenvalue decomposition of $G_{n}+\epsilon^{2}I=V\Sigma^{2}V^{*}$ we find that it suffices to show that
@@
$X(\mathbf{x})=w(\mathbf{x})\Sigma^{-1}V^{*}z(\mathbf{x})^{*}z(\mathbf{x})V\Sigma^{-1},$
where $\mathbf{x}$ is distributed according to $\mu$ and where $z(\mathbf{x})$ is defined as in Lemma 3. Then, $\Sigma^{-1}V^{*}G_{n,m}V\Sigma^{-1}$ is equal to the sum of $m$ i.i.d. copies of this random matrix, i.e.,
@@
Since $\operatorname{\mathop{{}\mathbb{E}}}(X_{j})=\Sigma^{-1}V^{*}G_{n}V\Sigma^{-1}$, we can use the matrix concentration result [10, Lemma 27]. Using Lemma 3, we bound
$\|X_{j}\|_{2}=w(\mathbf{x}_{j})z(\mathbf{x}_{j})(G_{n}+\epsilon^{2}I)^{-1}z(\mathbf{x}_{j})^{*}=w(\mathbf{x}_{j})k_{n}^{\epsilon}(\mathbf{x}_{j})$ $\leq K_{w,n}^{\epsilon},$
@@
where $K_{w,n}^{\epsilon}\coloneq\max_{\mathbf{x}\in X}\left(w(\mathbf{x})k_{n}^{\epsilon}(\mathbf{x})\right)$ and $I\in\mathbb{R}^{n\times n}$ is the identity matrix. Defining
@@
Appendix C Asymptotics of DPSWFs
In this section, we aim to motivate that the maxima $\|u_{i}\|_{L^{\infty}[-W,W]}$ associated to the one-dimensional Fourier extension frame (60) are $\mathcal{O}(\sqrt{n}\kern 1.0pt)$ as $n\to\infty$. In order to do so, observe that
$\|u_{i}\|_{L^{\infty}[-W,W]}=\|U_{i-1}(n,W)\|_{L^{\infty}[-W,W]}/\sqrt{\lambda_{i-1}(n,W)},$ (70)
where $U_{k}(n,W)$ and $\lambda_{k}(n,W)$ are the discrete prolate spheroidal wave functions (DPSWFs) and the associated eigenvalues analysed by Slepian in [46]. We recall and examine their asymptotic behaviour for the different regimes identified in [46, section 2.4-2.5]. Numerical evidence confirms that (70) is bounded by $\mathcal{O}(\sqrt{n})$, as is the case for the prolate spheroidal wave functions [14, Prop. 3.1], the continuous analogues of the DPSWFs.
Given the asymptotic expansions outlined in [46, section 2.4-2.5], we can derive that as $n\to\infty$:
@@ Appendix C Asymptotics of DPSWFs
    $\|u_{i}\|_{L^{\infty}[-W,W]}=\mathcal{O}(\sqrt{n}),$ (71)
@@
    $\|u_{i}\|_{L^{\infty}[-W,W]}\sim r(E\beta)\sqrt{\beta}\exp(\pi E\beta/4)\sqrt{\frac{3\pi(1+\exp(\pi b))}{\beta(1+\exp(\pi E\beta))\ln(n)}}\sqrt{n},$ (72)
    where $r$, $E$ and $\beta$ are defined by [46, eq. (52)-(54)],
@@
    $\|u_{i}\|_{L^{\infty}[-W,W]}\sim L_{2}^{-1/2}\pi(1-\cos(2\pi W)^{2})^{-1/4}\sqrt{n},$ (73)
    where $L_{2}$ is defined by [46, eq. (47)] with $A=-\cos(2\pi W)$ and $k=n-(i-1)$,
@@
    $\|u_{i}\|_{L^{\infty}[-W,W]}\sim\sqrt{\pi}\left(\frac{2}{1-\cos(2\pi W)}\right)^{1/4}\sqrt{n}.$ (74)
@@
$\|u_{i}\|_{L^{\infty}[-W,W]}$ $\leq\|u_{i}\|_{L^{\infty}[-1/2,1/2]}$
$=\|U_{i-1}(n,W)\|_{L^{\infty}[-1/2,1/2]}/\sqrt{\lambda_{i-1}(n,W)}$
@@
where we used the Nikolskii inequality [41, Theorem 1] with $\|U_{i-1}(n,W)\|_{L^{2}[-1/2,1/2]}=1$. Furthermore,
@@
using [46, eq. (60)]. Hence, we get $\|u_{i}\|_{L^{\infty}[-W,W]}=\mathcal{O}(\sqrt{n})$, where the proportionality factor is at most approximately $\sqrt{2}$.
For the case of $i\geq 2nW+1$, it follows from [43, Corollary 2] and the fact that the DPSWFs are even or odd that
$\|u_{i}\|_{L^{\infty}[-W,W]}=\frac{\|U_{i-1}(n,W)\|_{L^{\infty}[-W,W]}}{\sqrt{\lambda_{i-1}(n,W)}}=\frac{\lvert U_{i-1}(n,W;\omega=2\pi W)\rvert}{\sqrt{\lambda_{i-1}(n,W)}}.$
Hence, the maximum is achieved at the boundary. Furthermore, by combining [46, eq. (50) and (60)] for $U_{k}(n,W)$ and $\lambda_{k}(n,W)$ for $k=i-1=\lfloor 2nW+(b/\pi)\log(n)\rfloor$ with fixed $b$, one obtains
@@
Similarly, for $k=i-1=\lfloor 2nW(1+\epsilon)\rfloor$ with fixed $10<\epsilon<1/2W-$ 1 we derive from [46, eq. (42), (56), (57) and (63)] that
@@
while for $k=i-1=n-l$ with fixed $l\geq 1$ we derive from [46, eq. (13), (38) and (58)] that

### v2

@@ Appendix B Proof of Theorem 10
$k_{n}^{\epsilon}(\mathbf{x})=\phi(\mathbf{x})(G_{n}+\epsilon^{2}I)^{-1}\phi(\mathbf{x})^{*},$
where $\phi(\mathbf{x})=\begin{bmatrix}\phi_{1}(\mathbf{x})&\dots&\phi_{n^{\prime}}(\mathbf{x})\end{bmatrix}\in\mathbb{C}^{1\times n^{\prime}}$ and $G_{n}$ represents $\mathcal{T}_{n}^{*}\mathcal{T}_{n}$.
@@
$G_{n}$ is positive semidefinite and, hence, there exists an eigenvalue decomposition of the form $V\Sigma^{2}V^{*}$, where $V\in\mathbb{C}^{n^{\prime}\times n^{\prime}}$ is unitary and $\Sigma\in\mathbb{C}^{n^{\prime}\times n^{\prime}}$ is diagonal. Moreover, this implies that the synthesis operator $\mathcal{T}_{n}$ associated with $\Phi_{n}$ has a singular value decomposition of the form $\sum_{i=1}^{n}\sigma_{i}u_{i}\mathbf{v}_{i}^{*}$ where $\{u_{i}\}_{i=1}^{n}$ is an orthonormal basis for $V_{n}$, $\mathbf{v}_{i}$ are the first $n$ columns of $V$ and $\sigma_{i}=(\Sigma)_{i,i}$ for $i=1\dots n$. Therefore, using $\phi(\mathbf{x})=\sum_{i=1}^{n}\sigma_{i}u_{i}(\mathbf{x})\mathbf{v}_{i}^{*}$, we obtain
$\phi(\mathbf{x})(G_{n}+\epsilon^{2}I)^{-1}\phi(\mathbf{x})^{*}$ $=\left(\sum_{i=1}^{n}\sigma_{i}u_{i}(\mathbf{x})\mathbf{v}_{i}^{*}\right)V(\Sigma^{2}+\epsilon^{2}I)^{-1}V^{*}\left(\sum_{i=1}^{n}\sigma_{i}\overline{u_{i}(\mathbf{x})}\mathbf{v}_{i}\right)$
@@
The inequality on the left is exactly the $\Delta$-spectral approximation introduced in [12, eq. (2)] with $\Delta=1/2$. This allows us to reuse the method of proof of [12, Lemma 6]. Analogously, using the eigenvalue decomposition of $G_{n}+\epsilon^{2}I=V\Sigma^{2}V^{*}$ we find that it suffices to show that
@@
$X(\mathbf{x})=w(\mathbf{x})\Sigma^{-1}V^{*}\phi(\mathbf{x})^{*}\phi(\mathbf{x})V\Sigma^{-1},$
where $\mathbf{x}$ is distributed according to $\mu$ and where $\phi(\mathbf{x})$ is defined as in Lemma 4. Then, $\Sigma^{-1}V^{*}G_{n,m}V\Sigma^{-1}$ is equal to the sum of $m$ i.i.d. copies of this random matrix, i.e.,
@@
Since $\operatorname{\mathop{{}\mathbb{E}}}(X_{j})=\Sigma^{-1}V^{*}G_{n}V\Sigma^{-1}$, we can use the matrix concentration result [12, Lemma 27]. Using Lemma 4, we bound
$\|X_{j}\|_{2}=w(\mathbf{x}_{j})\phi(\mathbf{x}_{j})(G_{n}+\epsilon^{2}I)^{-1}\phi(\mathbf{x}_{j})^{*}=w(\mathbf{x}_{j})k_{n}^{\epsilon}(\mathbf{x}_{j})$ $\leq K_{w,n}^{\epsilon},$
@@
where $K_{w,n}^{\epsilon}\coloneq\max_{\mathbf{x}\in X}\left|w(\mathbf{x})k_{n}^{\epsilon}(\mathbf{x})\right|$ and $I\in\mathbb{R}^{n\times n}$ is the identity matrix. Defining
@@
Appendix C Proof of Theorem 11
Proof of Theorem 11.
For (58), consider
and
In order to obtain (59), we split the above as
$\|k_{n}^{\epsilon}\|_{\infty}$ $\leq\sup_{\mathbf{x}\in X}\left(\sum_{i=1}^{i_{1}}\lvert u_{i}(\mathbf{x})\rvert^{2}+\sum_{i=i_{1}+1}^{i_{2}}\lvert u_{i}(\mathbf{x})\rvert^{2}+\operatorname{S_{\text{tail}}}\max_{i_{2}<i\leq n}\lvert u_{i}(\mathbf{x})\rvert^{2}\right)$
$\leq\sup_{\mathbf{x}\in X}\sum_{i=1}^{i_{1}}\lvert u_{i}(\mathbf{x})\rvert^{2}+((i_{2}-i_{1})+\operatorname{S_{\text{tail}}})\max_{i_{1}<i\leq n}\|u_{i}\|_{\infty}^{2}.$
$\sum_{i=1}^{i_{1}}\lvert u_{i}(\mathbf{x})\rvert^{2}=\|\phi(\mathbf{x})VD^{-1}\|_{2}^{2}\leq\|\phi(\mathbf{x})\|_{2}^{2}\|V\|_{2}^{2}\|D^{-1}\|_{2}^{2}=\frac{\|\phi(\mathbf{x})\|_{2}^{2}}{\sigma_{i_{1}}^{2}}\leq\frac{\|\phi(\mathbf{x})\|_{2}^{2}}{1-\epsilon^{2}},$
where $\phi(\mathbf{x})=\begin{bmatrix}\phi_{1,n}(\mathbf{x})&\dots&\phi_{n,n}(\mathbf{x})\end{bmatrix}$, $V\in\mathbb{C}^{n\times i_{1}}$ is a matrix with orthogonal columns and $D\in\mathbb{C}^{i_{1}\times i_{1}}$ is a diagonal matrix with $(D)_{(i,i)}=\sigma_{i}$. The final result follows from
$\sup_{\mathbf{x}\in X}\frac{\|\phi(\mathbf{x})\|_{2}^{2}}{1-\epsilon^{2}}\leq\frac{1}{1-\epsilon^{2}}\sup_{\mathbf{x}\in X}\sum_{i=1}^{n}\lvert\phi_{i,n}(\mathbf{x})\rvert^{2}\leq\frac{1}{1-\epsilon^{2}}\sum_{i=1}^{n}\|\phi_{i,n}\|_{\infty}^{2}\leq\frac{n}{1-\epsilon^{2}}.$
Appendix D Asymptotics of DPSWFs
In this section, we aim to motivate that the maxima $\|u_{i}\|_{\infty}$ associated to the one-dimensional Fourier extension frame (64) are $\mathcal{O}(\sqrt{n}\kern 1.0pt)$ as $n\to\infty$. In order to do so, observe that
$\|u_{i}\|_{\infty}=\|U_{i-1}(n,W)\|_{\infty}/\sqrt{\lambda_{i-1}(n,W)},$ (70)
where $U_{k}(n,W)$ and $\lambda_{k}(n,W)$ are the discrete prolate spheroidal wave functions (DPSWFs) and the associated eigenvalues analysed by Slepian in [50]. We recall and examine their asymptotic behaviour for the different regimes identified in [50, section 2.4-2.5]. Numerical evidence confirms that (70) is bounded by $\mathcal{O}(\sqrt{n})$, as is the case for the prolate spheroidal wave functions [16, Prop. 3.1], the continuous analogues of the DPSWFs.
Given the asymptotic expansions outlined in [50, section 2.4-2.5], we can derive that as $n\to\infty$:
@@ Appendix D Asymptotics of DPSWFs
    $\|u_{i}\|_{\infty}=\mathcal{O}(\sqrt{n}),$ (71)
@@
    $\|u_{i}\|_{\infty}\sim r(E\beta)\sqrt{\beta}\exp(\pi E\beta/4)\sqrt{\frac{3\pi(1+\exp(\pi b))}{\beta(1+\exp(\pi E\beta))\ln(n)}}\sqrt{n},$ (72)
    where $r$, $E$ and $\beta$ are defined by [50, eq. (52)-(54)],
@@
    $\|u_{i}\|_{\infty}\sim L_{2}^{-1/2}\pi(1-\cos(2\pi W)^{2})^{-1/4}\sqrt{n},$ (73)
    where $L_{2}$ is defined by [50, eq. (47)] with $A=-\cos(2\pi W)$ and $k=n-(i-1)$,
@@
    $\|u_{i}\|_{\infty}\sim\sqrt{\pi}\left(\frac{2}{1-\cos(2\pi W)}\right)^{1/4}\sqrt{n}.$ (74)
@@
$\|u_{i}\|_{\infty,[-W,W]}$ $\leq\|u_{i}\|_{\infty,[-1/2,1/2]}$
$=\|U_{i-1}(n,W)\|_{\infty,[-1/2,1/2]}/\sqrt{\lambda_{i-1}(n,W)}$
@@
where we used the Nikolskii inequality [43, Theorem 1] with $\|U_{i-1}(n,W)\|_{\infty,[-1/2,1/2]}=1$. Furthermore,
@@
using [50, eq. (60)]. Hence, we get $\|u_{i}\|_{\infty,[-W,W]}=\mathcal{O}(\sqrt{n})$, where the proportionality factor is at most approximately $\sqrt{2}$.
For the case of $i\geq 2nW+1$, it follows from [47, Corollary 2] and the fact that the DPSWFs are even or odd that
$\|u_{i}\|_{\infty}=\frac{\|U_{i-1}(n,W)\|_{\infty}}{\sqrt{\lambda_{i-1}(n,W)}}=\frac{\lvert U_{i-1}(n,W;\omega=2\pi W)\rvert}{\sqrt{\lambda_{i-1}(n,W)}}.$
Hence, the maximum is achieved at the boundary. Furthermore, by combining [50, eq. (50) and (60)] for $U_{k}(n,W)$ and $\lambda_{k}(n,W)$ for $k=i-1=\lfloor 2nW+(b/\pi)\log(n)\rfloor$ with fixed $b$, one obtains
@@
Similarly, for $k=i-1=\lfloor 2nW(1+\epsilon)\rfloor$ with fixed $10<\epsilon<1/2W-$ 1 we derive from [50, eq. (42), (56), (57) and (63)] that
@@
while for $k=i-1=n-l$ with fixed $l\geq 1$ we derive from [50, eq. (13), (38) and (58)] that